Circular No. 8606

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 S2 (SKIFF) B. Skiff, Lowell Observatory, reports his discovery of a comet (discovery

observation tabulated below) on CCD images taken with the LONEOS 0.59m Schmidt telescope, the object showing a moderately condensed 6′′ coma and a fan-shaped tail extending perhaps 15′′ toward the southwest (in the direction of motion). A 5-min exposure taken by Skiff with the Perkins 1.8m telescope (+ PRISM camera + ‘VR’ filter) shows the coma diameter as ∼ 5′′ and the tail extending at least 36′′ in p.a. ∼ 240^{o}. Following posting on the ‘NEO Confirmation Page’, J. Young reports that six stacked CCD images taken on Sept. 30.23–30.28 UT with the 0.6-m Cassegrain reflector at Table Mountain show the comet to have a 4′′ coma with a 12′′ tail in p.a. 260^{o}, and E. J. Christensen reports that four 45-s co-added frames taken with the 1.5-m Catalina reflector on Sept. 30.3 show an 8′′ coma and a 20′′ tail in p.a. 230^{o}–240^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.29.27649  0^{h}55^{m}25.^{s}73  +3^{o}48′39′′.9  19.1

The available astrometry (including Sept. 25 prediscovery Spacewatch observations), the following very indeterminate parabolic orbital elements, and an ephemeris appear on MPEC 2005-S78.

T = 2010 Feb. 12.081 TT  ω = 306.^{o}165
Ω = 201.227 }2000.0
q = 3.29967 AU  i = 166.711

SUPERNOVA 2005ee IN PGC 73054 M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics, report that a spectrogram (range 340–740 nm) of SN 2005ee (cf. IAUC 8603), obtained by N. Martimbeau on Sept. 25.32 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-II supernova. The spectrum consists of a flat continuum, a P-Cyg profile of Hα, and absorption troughs of Hβ and other lines. Adopting the recession velocity of 9730 km/s for the host galaxy, as measured from Hα emission of an H II region in the host galaxy, the supernova expansion velocity derived from the minimum of the Hβ line is ∼ 11000 km/s.

2005 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green